OpenStudy (anonymous):

Solve the formula for the variable y. z = 4x + 2y

OpenStudy (anonymous):

z=4x+2y -4x from each side z-4x = 2y divide each side by 2 \[\frac{ z-4x }{ 2 }\]

OpenStudy (anonymous):

that equals y

OpenStudy (anonymous):

y=z/2 - 2x

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

thanks for the help guys !

OpenStudy (anonymous):

z - 4x = 2y \[y = \frac{ z-4x }{ 2 }\] y = \[y = \frac{ 1 }{ 2 }z - 2\]

OpenStudy (anonymous):

z = 4x + 2y, solve for Y Move the Z and X variables to the same side of the equation by subtracting 4x from both sides, that leaves you with: z - 4x = 2y To isolate the y variable by itself on that side of the equation, divide both sides by 2, so: (z-4x)/2 = y This can be simplified to: (z/2) - 2x = y